OpenStudy (anonymous):

(2/(x-2))=(3/(x+5))+(10/(x+5)(x-2))

OpenStudy (anonymous):

Use P.E.M.D.A.S, its pretty simple. Parentheses Exponents Multiplication Division Addition Subtraction. Go through and start solving in that order. I got x = 6.

OpenStudy (anonymous):

@Dihas You do realize that these are fractions right?

OpenStudy (anonymous):

Honestly you should have provided more information on the type of answer you were looking for. Posting just that line makes it hard to tell what it is, and what you're trying to get out of it.

OpenStudy (anonymous):

@Dihas I don't know what other information there is to say. All you have to do is find x. By using parenthesis, it's clear that I was using fractions. Is it really necessary to say: "Find the x from the equation of fractions"? I don't think so. If you don't know it, it's safe to just not answer

OpenStudy (anonymous):

\[\frac{ 2 }{ (x-2) }=\frac{ 3 }{ (x+5) }+\frac{ 10 }{ (x+5)(x-2) }\] right? first cross multiply the right-hand side, \[\frac{ 2 }{ (x-2) }=\frac{ 3 { (x+5)(x-2) + { 10 { (x+5) } } } }{ (x+5)(x+5)(x-2) }\]then cancel out "x+5" since all the three terms on the right-hand side has it. NOTE: in the denominator, you only cancel 1 "x+5". \[\frac{ 2 }{ (x-2) }=\frac{ 3 {(x-2) + { 10 } } }{ (x+5)(x-2) }\] after that, multiply both side by "x-2" so that the denominator on the left-hand side will be cancelled out. \[\ 2 =\frac{ 3 {(x-2)(x-2) + { 10 (x-2) } } }{ (x+5)(x-2) }\]again, cancel out "x-2" on the right hand side. \[\ 2 =\frac{ 3 {(x-2) + { 10 } } }{ (x+5) }\]cross-multiply, \[\ 2{ (x+5) } ={ 3 {(x-2) + { 10 } } }\]apply distributive property, \[\ 2x+10 = 3x-6+10\]solve for x, \[\ 2x+10 = 3x+4\]\[\ 10-4 = 3x-2x\]\[\ 6 = x\]\[\ x = 6\] that's your answer. oh by the way, please show some respect to people that responds to your problem.